# MambaLRP: Explaining Selective State Space Sequence Models

### v1

@@ A.3 Derivations for multiplicative gate
Appendix B Explicit propagation rules for MambaLRP
@@ C.1 Models and datasets
For the NLP experiments, we fine-tuned all parameters of the pretrained Mamba-130M and Mamba-1.4B models [^2] https://github.com/state-spaces/mamba on four text classification datasets: SST-2, SNLI, Medical BIOS, and Emotion. The data statistics can be seen in Table C.1. For the vision experiments, we used the pretrained Vim-S model [^3] https://github.com/hustvl/Vim, trained on the ImageNet dataset.
@@
During training, we used a batch size of 32. To train the Mamba-1.4B model on the SNLI dataset, a batch size of 64 is used. We employed the {EleutherAI/gpt-neox-20b} [^4] https://github.com/EleutherAI/gpt-neox tokenizer. The models’ parameters were optimized using AdamW optimizer with a learning rate set at $57e-$ 5. Additionally, we used a linear learning rate scheduler with an initial factor of 0.5. All models were trained for a maximum of 10 epochs, with an early stopping mechanism in place. The top-1 accuracies of the models on each dataset are detailed in Table C.1.
Table 4: The accuracies of Mamba-130M and Mamba-1.4B models on the validation sets of four text classification datasets.
Dataset Mamba-130M Mamba-1.4B
SST-2 91.97 94.15
Med-BIOS 89.10 90.30
Emotion 93.45 93.65
SNLI 89.57 91.05
Table 5: Data statistics.
@@ C.2.1 LRP composites for Vision Mamba
Table 6: Finding the best LRP composite for Vision Mamba. The layers in which the generalized LRP-$\gamma$ rule is applied are represented with LRP-$\gamma$ and the ones in which the basic LRP rule, i.e. LRP-0, is used are represented with LRP-0.
@@
LRP-$\gamma$ LRP-$\gamma$ LRP-$\gamma$ 1.7173
LRP-$\gamma$ LRP-$\gamma$ LRP-0 1.7315
LRP-0 LRP-$\gamma$ LRP-$\gamma$ 1.7736
LRP-0 LRP-$\gamma$ LRP-0 1.7824
LRP-$\gamma$ LRP-0 LRP-0 1.8588
LRP-$\gamma$ LRP-0 LRP-$\gamma$ 1.8646
LRP-0 LRP-0 LRP-0 1.8852
LRP-0 LRP-0 LRP-$\gamma$ 1.9106
@@ C.3 Further details of other explanation methods
Table 7: Hyperparameters of other explanation methods. The parameters $\mu$ and $\sigma$ represent the mean and standard deviation of noise, respectively, while the parameter $m$ denotes the sample size.
@@
In the vision experiments, we used the original implementations  [^5] https://github.com/AmeenAli/HiddenMambaAttn/ of the AttnRoll and G×$\times$×AttnRoll methods, provided to explain the Vim-S model. Given the unavailability of code for adapting these approaches to the language models, namely Mamba-130M and Mamba-1.4B, we have developed our own implementation. In the vision case, the authors obtain the final relevance map by extracting the row associated with the CLS token in the attention matrix. However, since our language models lack a CLS token, we get the final relevance map from the row associated with the last token in the attention matrix. This is because predictions are based on the last state in these models. For the gradient-based methods, we use the implementations available in the Captum library [^6] https://captum.ai/.
C.4 Additional quantitative results
The results of the insertion experiment are shown in Table 8. Similar to the flipping results, MambaLRP outperforms all baseline methods for all the datasets and models. In the majority of the models and datasets, the G×$\times$×AttnRoll method has shown better performance compared to the pure gradient-based approaches.
Table 8: Faithfulness score $\Delta{A}^{I}\uparrow$. A higher score indicates more faithful explanations.
Methods SST-2 Med-BIOS SNLI Emotion ImageNet
Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Mamba 130M Mamba 1.4B Vim-S
Random -0.024 0.009 0.022 0.004 -0.003 0.001 0.019 0.021 0.000
GI [54] 0.074 -0.108 0.200 -0.649 -0.043 -0.040 -0.839 -0.533 0.041
SmoothGrad [58] 1.399 -0.405 1.655 -2.295 0.497 -0.709 1.904 -1.982 0.121
IG [55] 0.880 0.223 1.312 1.065 0.465 0.224 1.909 2.181 0.328
AttnRoll [4] 0.704 0.554 2.265 1.105 0.257 0.375 0.450 1.656 0.714
G $\times$ AttnRoll [4] 1.238 0.751 3.175 3.032 0.534 0.564 2.091 4.963 0.794
LRP (LN-rule, [3]) 0.916 1.083 2.235 3.471 0.718 0.664 3.208 5.443 0.648
MambaLRP (ours) 2.038 1.379 3.933 4.250 1.024 0.917 3.640 5.650 1.878
@@ C.5 Further ablation experiments
In Section 4.3, we proposed several strategies to mitigate conservation violation in the the Mamba block’s multiplicative gate. In this experiment, we evaluate the proposed approaches. As can be seen in Table 9, detaching the multiplicative gate leads to lower faithfulness scores compared to the half-relevance propagation approach. To retain conservation, an alternative approach is to detach the SSM’s output, which limits capturing long-range dependencies, a task for which this branch is designed for. Detaching it may result in a loss of valuable information used by the model to make predictions.
Table 9: Comparing the proposed strategies for managing the Mamba block’s multiplicative gate.
@@
Detaching multiplicative gate 1.577 1.387
Half-relevance propagation 1.978 1.878
@@ C.6.1 Natural language processing
In the following figures, we represent explanations produced by MambaLRP and other baseline methods to interpret the Mamba-130M models trained on various datasets. In the visualizations, shades of red represent words that positively influence the model’s prediction. Conversely, shades of blue reflect negative contributions. The heatmaps of the AttnRoll and G×$\times$×AttnRolll methods are constrained to non-negative values.
Figure 8: Explanations generated by different explanation methods for a sentence of the SNLI validation set. This sentence belongs to the ‘entailment’ class.
Figure 9: Explanations generated by different explanation methods for a sentence of the SNLI validation set. This sentence belongs to the ‘contradiction’ class.
Figure 10: Explanations generated by different explanation methods for a sentence of the Emotion validation set. This sentence belongs to the ‘joy’ class.
Figure 11: Explanations generated by different explanation methods for a sentence of the Medical BIOS validation set. This sentence belongs to the ‘nurse’ class.
@@ C.6.2 Computer vision
Figure 12: Explanations produced by different explanation methods for images of the ImageNet dataset. Explanations produced by AttnRoll and G×$\times$×AttnRoll are limited to non-negative values, whereas those generated by gradient-based techniques and MambaLRP includes both positive and negative contributions.
@@ C.7 Additional use case results
For the needle-in-a-haystack experiment in Section 6, we use a synthetic dataset  [^7] https://huggingface.co/datasets/lvwerra/needle-llama3-16x512. In this dataset, a single passkey (the ‘needle’) is inserted at different locations within a collection of repeated noise sentences (the ‘haystack’), as described in [33]. The dataset is composed of sequences with different context lengths. In our experiment, we use sequences with context lengths of 512, 1024, and 2048. We use a Mamba-2.8B model  [^8] https://huggingface.co/clibrain/mamba-2.8b-chat-no_robots, which is finetuned on the No Robots dataset [46] using a context length of 2048. Then, we prompt the model to extract the passkey hidden among irrelevant text by completing the phrase "The passkey is  ".
@@
Retrieval accuracy is a metric, which is commonly used in the needle-in-a-haystack experiment to analyze the model’s performance. The synthetic dataset used for this experiment can be designed to include misleading information, which may cause the model to generate the correct passkey based on incorrect evidence. In such cases, simply evaluating the retrieval accuracy may be insufficient. This issue can also arise when dealing with more realistic haystacks. Therefore, we introduced explanation-based retrieval accuracy (XRA) in Section 7. MambaLRP and the XRA metric designed upon it can help to better examine the evidence the model relies on to retrieve the needle. In our experiment, we set the value of K to 2. This is because LRP usually identifies the token immediately preceding the generated token as the most important one and the evidence used for the passkey retrieval is usually the second most important token.
@@
The sample in Fig. 13 represents such scenario. In this case, the next token generated by the model is the second part of the correct passkey (300). However, the model has incorrectly focused on the number 300 in the phrase "Pass the key to room 6300" to generate this token. Simply looking at the retrieved token might suggest that the model successfully retrieved the correct information. However, examining the MambaLRP’s explanation heatmaps provides deeper insights into the model’s behavior. This helps us to debug the model more effectively and design better tests to analyze its capabilities.
@@
Figure 13: Detecting Clever-Hans effect in the needle-in-a-haystack test. Given the 2K context length in this example, visualizing the entire text could be confusing. Therefore, we have removed most of the haystack from the visualization. In this example, the model has generated the correct passkey but the generation is not based on truly relevant information in the text.
C.8 Runtime comparison
@@
Table 10: Runtime comparison. The time needed for each baseline method to generate its explanations. The times, measured in seconds, are averaged over the samples from the Medical BIOS dataset. The model used in this experiment is Mamba-130M without using fast CUDA kernels.
@@
Table 11: Runtime comparison. The time needed for each baseline method to generate its explanations. The times, measured in seconds, are averaged over the samples from the Medical BIOS dataset. The model used in this experiment is Mamba-130M using fast CUDA kernels.

### v3

@@ A.3 Derivations for multiplicative gate
Приложение B Explicit propagation rules for MambaLRP
@@ C.1 Models and datasets
For the NLP experiments, we fine-tuned all parameters of the pretrained Mamba-130M, Mamba-1.4B, and Mamba-2.8B models [^2] https://github.com/state-spaces/mamba on four text classification datasets: SST-2, SNLI, Medical BIOS, and Emotion. The data statistics can be seen in Table C.1. For the vision experiments, we used the pretrained Vim-S model [^3] https://github.com/hustvl/Vim, trained on the ImageNet dataset.
@@
During training, we used a batch size of 32. To train the Mamba-1.4B and Mamba-2.8B models on the SNLI dataset, a batch size of 64 is used. We employed the {EleutherAI/gpt-neox-20b} [^4] https://github.com/EleutherAI/gpt-neox tokenizer. The models’ parameters were optimized using AdamW optimizer with a learning rate set at $57e-$ 5. Additionally, we used a linear learning rate scheduler with an initial factor of 0.5. All models were trained for a maximum of 10 epochs. We employed early stopping and ended training as soon as the validation loss ceased to improve. The top-1 accuracies of the models on each dataset are detailed in Table C.1.
Таблица 4: The accuracies of Mamba-130M, Mamba-1.4B, and Mamba-2.8B models on the validation sets of four text classification datasets.
Dataset 130M 1.4B 2.8B
SST-2 91.97 94.15 94.26
Med-BIOS 89.10 90.30 90.50
Emotion 93.45 93.65 94.10
SNLI 89.57 91.05 91.14
Таблица 5: Data statistics.
@@ C.2.1 LRP composites for Vision Mamba
Таблица 6: Finding the best LRP composite for Vision Mamba. The layers in which the generalized LRP-$\gamma$ rule is applied are represented with LRP-$\gamma$ and the ones in which the basic LRP rule, i.e. LRP-0, is used are represented with LRP-0.
@@
LRP-$\gamma$ LRP-$\gamma$ LRP-$\gamma$ 4.196
LRP-$\gamma$ LRP-$\gamma$ LRP-0 4.218
LRP-0 LRP-$\gamma$ LRP-$\gamma$ 4.283
LRP-0 LRP-$\gamma$ LRP-0 4.336
LRP-$\gamma$ LRP-0 LRP-$\gamma$ 4.597
LRP-$\gamma$ LRP-0 LRP-0 4.599
LRP-0 LRP-0 LRP-0 4.684
LRP-0 LRP-0 LRP-$\gamma$ 4.715
@@ C.3 Further details of other explanation methods
Таблица 7: Hyperparameters of other explanation methods. The parameters $\mu$ and $\sigma$ represent the mean and standard deviation of noise, respectively, while the parameter $m$ denotes the sample size.
@@
In the vision experiments, we used the original implementations  [^5] https://github.com/AmeenAli/HiddenMambaAttn/ of the AttnRoll and MambaAttr methods, provided to explain the Vim-S model. Given the unavailability of code for adapting these approaches to the language models, namely Mamba-130M, Mamba-1.4B, and Mamba-2.8B, we have developed our own implementation. In the vision case, the authors obtain the final relevance map by extracting the row associated with the CLS token in the attention matrix. However, since our language models lack a CLS token, we get the final relevance map from the row associated with the last token in the attention matrix. This is because predictions are based on the last state in these models. For the gradient-based methods, we use the implementations available in the Captum library [^6] https://captum.ai/.
C.4 Further details on evaluation metrics
As mentioned in Section 5.3, the flipping and insertion metrics can be used to evaluate the quality of the generated explanations. It is important to note that starting with unperturbed images and gradually applying perturbations until fully perturbed images are obtained yields identical results for both the flipping and insertion metrics. Therefore, we have only reported the results for the flipping experiment.
In our flipping evaluations, we calculate the area under the curve (AUC) by starting from full images and progressively masking the pixels with zeros until we reach completely masked images. The perturbation steps are defined using np.linspace(0, 1, 11) . In our vision experiments, Images are normalized using the ImageNet mean and standard deviation, and the explanations are generated and evaluated for the predicted class. Unlike [4], which tracks changes in the model’s top-1 accuracy, we monitor changes in the output logit of the predicted class.
@@ C.5 Further ablation experiments
In Section 4.3, we proposed several strategies to mitigate conservation violation in the Mamba block’s multiplicative gate. In this experiment, we evaluate the proposed approaches. As can be seen in Table 8, detaching the multiplicative gate $z_{B}$ leads to lower faithfulness scores compared to the half-relevance propagation approach. To retain conservation, an alternative approach is to detach the SSM’s output $z_{A}$, which limits capturing long-range dependencies, a task for which this branch is designed for. Detaching it may result in a loss of valuable information used by the model to make predictions. Therefore, this approach is not considered in Table 8.
Таблица 8: Comparing the proposed strategies for managing the Mamba block’s multiplicative gate.
@@
Detaching the multiplicative gate $z_{B}$ 1.577 3.592
Half-relevance propagation 1.978 4.715
@@ C.6.1 Natural language processing
In the following figures, we represent explanations produced by MambaLRP and other baseline methods to interpret the Mamba-130M models trained on various datasets. In the visualizations, shades of red represent words that positively influence the model’s prediction. Conversely, shades of blue reflect negative contributions. The heatmaps of the AttnRoll and MambaAttr methods are constrained to non-negative values.
Рис. 8: Explanations generated by different explanation methods for a sentence of the SNLI validation set. This sentence belongs to the ‘entailment’ class.
Рис. 9: Explanations generated by different explanation methods for a sentence of the SNLI validation set. This sentence belongs to the ‘contradiction’ class.
Рис. 10: Explanations generated by different explanation methods for a sentence of the Emotion validation set. This sentence belongs to the ‘joy’ class.
Рис. 11: Explanations generated by different explanation methods for a sentence of the Medical BIOS validation set. This sentence belongs to the ‘nurse’ class.
@@ C.6.2 Computer vision
Рис. 12: Explanations produced by different explanation methods for images of the ImageNet dataset. Explanations produced by AttnRoll and MambaAttr are limited to non-negative values, whereas those generated by gradient-based techniques and MambaLRP includes both positive and negative contributions.
@@ C.7 Additional use case results
For the needle-in-a-haystack experiment in Section 6, we use a synthetic dataset  [^7] https://huggingface.co/datasets/lvwerra/needle-llama3-16x512. In this dataset, a single passkey (the ‘needle’) is inserted at different locations within a collection of repeated noise sentences (the ‘haystack’), as described in [37]. The dataset is composed of sequences with different context lengths. In our experiment, we use sequences with context lengths of 512, 1024, and 2048. We restrict the maximum context length to 2048 tokens to align with the model’s training configuration, as this experiment is not designed to evaluate extrapolation beyond this limit. The goal is to focus on certain limitations of the retrieval accuracy metric and and the solutions provided by MambaLRP. We use a Mamba-2.8B model  [^8] https://huggingface.co/clibrain/mamba-2.8b-chat-no_robots, which is finetuned on the No Robots dataset [55] using a context length of 2048. Then, we prompt the model to extract the passkey hidden among irrelevant text by completing the phrase ‘‘The passkey is  ".
@@
Retrieval accuracy is a metric, which is commonly used in the needle-in-a-haystack experiment to analyze the model’s performance. The synthetic dataset used for this experiment can be designed to include misleading information, which may cause the model to generate the correct passkey based on incorrect evidence. In such cases, simply evaluating the retrieval accuracy may be insufficient. This issue can also arise when dealing with more realistic haystacks. Therefore, we introduced explanation-based retrieval accuracy (XRA) in Section 6. MambaLRP and the XRA metric designed upon it can help to better examine the evidence the model relies on to retrieve the needle. In this approach, we first identify the positions of the top-K relevant tokens by MambaLRP, and then, calculate the accuracy by comparing those positions to the needle’s position. We set the value of K to 2. This is because MambaLRP identifies the token immediately preceding the generated token as the most important one in most of the examples and the evidence used for the passkey retrieval is usually the second most important token.
@@
The sample in Fig. 13 represents a scenario where our XRA approach proves valuable. In this case, the next token generated by the model is the second part of the correct passkey (300). However, the model has incorrectly focused on the number 300 in the phrase ‘‘Pass the key to room 6300"to generate this token. Simply looking at the retrieved token might suggest that the model successfully retrieved the correct information. However, examining the MambaLRP’s explanation heatmaps provides deeper insights into the model’s behavior. This helps us to debug the model more effectively and design better tests to analyze its capabilities.
@@
Рис. 13: Detecting Clever-Hans effect in the needle-in-a-haystack test. Given the 2K context length in this example, visualizing the entire text could be confusing. Therefore, we have removed most of the haystack from the visualization. In this example, the model has generated the correct passkey but the generation is not based on truly relevant information in the text.
C.8 Long-range dependeny comparison to Transformers
To explore the capabilities of different model architectures in handling long-range dependencies, we performed a direct comparison between Mamba and state-of-the-art Transformers (Llama-2999https://huggingface.co/meta-llama/Llama-2-7b-hf and Llama-3101010https://huggingface.co/meta-llama/Meta-Llama-3-8B), focusing on their performance with inputs exceeding typical context lengths.
For Llama-2 and Llama-3, we extract attributions using LRP for Transformers [3]. As in the Mamba experiment, we generate 10 additional tokens from the HotPotQA dataset input and explain the prediction for each generated token. The results are shown in Table 9. For Llama-2, which was trained with a context length of 4096, the generated text becomes increasingly less sensible and repetitive for contexts longer than 4k, a limitation noted also in [19, 5]. When analyzing the histogram distribution over tokens considered relevant to predict the next token, it appears that Llama-2 uses information more uniformly across the entire context and identifies more relevant long-range dependent tokens compared to Llama-3 and Mamba. However, as presented in Table 9, its output becomes nonsensical for context lengths exceeding 4K tokens, characterized by the use of rare vocabulary and repeated tokens. Thus, the identified relevant tokens are mostly non-semantic such as new line token ‘$<$0x0A$>$’ in Llama-2 and beginning of sentence token ‘$<$s$>$’ found at the start of the context paragraphs. For Llama-3 and Mamba, the attributions can identify meaningful relevant tokens. When directly compared, Llama-3 uses information from more intermediate mid-range dependencies than Mamba, though both favor tokens close to the end of the input as relevant. Given Llama-3’s much larger size (8B) compared to Mamba (130M) and their different training settings, this analysis supports that Mamba indeed uses long-range information. We also find that this ability is not exclusive to SSMs and can in principle also be achieved by Transformer models.
To what extent these findings depend on the amount of training data and model complexity remains an open research question. Our investigation of long-range dependencies in recent sequence generation models highlights the value of faithful attribution methods like MambaLRP in examining the capabilities and mechanisms utilized by models during generation.
id length Llama-2 Llama-3 Mamba
3 1k Summary <0x0A> The  genus  D ict y os per ma  is  a Summary Ċ The genus Dict y os per ma is a mon Summary Ċ The species is a member of the genus Ap oll
41 4k Summary <0x0A> The  Ohio ,  Ohio ,  Ohio ,  Ohio ,  Ohio Summary Ċ The following is a summary of Finn ’s head coaching Summary Ċ The following is a summary of the history of the
109 8k Summary <0x0A> The Љ Љ Љ Љ Љ Љ Љ Љ Љ Summary Ċ The mall is anchored by Hudson ’s Bay , Walmart Summary Ċ The mall is located in the heart of the city
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Таблица 9: Long-Range dependency experiment, comparing Transformers and Mamba for different context lengths.
C.9 Runtime comparison
@@
Таблица 10: Runtime comparison. The time needed for each baseline method to generate its explanations. The times, measured in seconds, are averaged over the samples from the Medical BIOS dataset. The model used in this experiment is Mamba-130M without using fast CUDA kernels.
@@
Таблица 11: Runtime comparison. The time needed for each baseline method to generate its explanations. The times, measured in seconds, are averaged over the samples from the Medical BIOS dataset. The model used in this experiment is Mamba-130M using fast CUDA kernels.